Status Report

NASA ISS On-Orbit Status 29 October 2008

By SpaceRef Editor
October 30, 2008
Filed under , , ,
NASA ISS On-Orbit Status 29 October 2008

All ISS systems continue to function nominally, except those noted previously or below.

CDR Fincke set up the camcorder equipment for taking video of Chamitoff and himself performing their first session of the general U.S. MedOps PFE (Periodic Fitness Evaluation) program on the CEVIS (Cycle Ergometer with Vibration Isolation), filmed via VTR (Video Tape Recorder). [The footage was downlinked afterwards for biomechanical evaluation of the exercising crewmember and assessment of the on-orbit setup of equipment during data collection and hardware status.]

The CDR & FE-2 then completed the PFE protocol, a monthly 1.5-hr. procedure which checks up on blood pressure and electrocardiogram (ECG) during programmed exercise on the CEVIS (Cycle Ergometer with Vibration Isolation) in the Lab. Readings were taken with the BP/ECG (blood pressure/electrocardiograph) and the HRM (heart rate monitor) watch with its radio transmitter, with Chamitoff first assisting as CMO (Crew Medical Officer), then being assisted by Fincke in his turn. The video equipment was later stowed and MCC-Houston informed that the video could be downlinked. [BP/ECG provides automated noninvasive systolic and diastolic blood pressure measurements while also monitoring and displaying accurate heart rates on a continual basis at rest and during exercise.]

FE-1 Lonchakov had ~4.5h to perform periodic replacement maintenance on the Russian thermal control system’s heating loop #2 (SOTR KOB-2), removing two units (BS) containing electric pumps (ENA) of the replaceable pump panel 4SPN2 and exchanging them with spares from stowage. Afterwards, Yuri also tested the 4SPN2 pump panel to ascertain its functionality. Today’s maintenance leaves one remaining spare on board. Finally, the KOB-2 was turned on by Russian ground control. [The two BS units of 4SPN2 were exchanged the last time on 8/22. Each of the two SOTR KOB loops has two redundant pump panels with two redundant pumps each. While in the early years of Mir and ISS the pumps were integral to the SPN panels, the current design allows them to be replaced without requiring an entire new SPN block.]

Gregory Chamitoff continued his support of the EarthKAM (Earth Knowledge Acquired by Middle Schools) program, today swapping the 60mm lens with a 180mm lens for the students. [There are 65 schools and 4,768 students participating in this week’s session. EarthKAM allows middle school students to program a digital camera onboard the ISS to photograph a variety of geographical targets for study in the classroom.]

Mike & Greg continued to prepare for the STS-126/ULF-2 mission by prepacking hardware for return in the MPLM (Multi-Purpose Logistics Module).

CDR Fincke reconfigured the toolbox facility in Node-1, installing new drawers and rearranging old drawers. [Mike prepared a common crew tool kit for use in Node-2, Columbus, Kibo JPM (JEM Pressurized Module) and the MPLM coming up with ULF-2, added a few tools to the ISS Leak Kit, removed/stowed tools from old drawers, and stowed old empty drawers for return on ULF-2.]

FE-2 Chamitoff replaced the PS1 & PS2 dust filters in the FGB with new spares.

Lonchakov collected the periodic readings of potentially harmful atmospheric contaminants in the SM (Service Module), using the CMS (Countermeasure System), a component of the SKDS GANK-4M Real-Time Harmful Contaminant Gas Analyzer suite, today using preprogrammed microchips to measure for o-Xylol (1,2-Dimethylbenzol, C8H10), Methyl-Mercaptan (Methanethiol, CH4S), and Sulphur Dioxide (SO2).

The crew reported that the aft Airlock GLA (General Luminaire Assembly) was dim. Troubleshooting was performed on the light and power removed to protect subsequent failure of the Baseplate Ballast Assembly (BBA). Engineering is recommending replacing the Lamp Housing Assembly (LHA).

The SSRMS (Space Station Remote Manipulator System) was successfully repositioned today in preparation for the ESP-3 (External Stowage Platform 3) relocation. The PDGF-3 (Power & Data Grapple Fixture #3) was released and the arm was maneuvered and grappled PDGF-4. A base change was performed from PDGF-3 to PDGF-4. The Tip LEE (Latching End Effector) on PDGF-1 will be released tomorrow and the arm will be maneuvered to the ESP-3 park position. Additionally, the demate microswitch on the MBS (Mobile Base System) redundant string was confirmed healthy today. [This microswitch failed to indicate demate during the MSS MCAS (Mobile Service System Common Attach System) checkout on 10/7.]

Yuri Lonchakov had an hour to himself for general orientation (station familiarization & acclimatization) as is standard daily rule for fresh crewmembers for the first two weeks after starting station residence, if she/he chooses to take it.

The crew executed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-1, FE-2) and VELO bike with bungee cord load trainer (FE-1). [Maximum speed for TVIS in motorized mode is not to exceed 16.1 km/h. Fincke removed his SPDs (Subject Positioning Devices) which were required for the first seven exercise sessions for him for safety.]

After a successful second day of the HTV PROX (H-II Transfer Vehicle/Prox System) Checkout, HTVCC (HTV Control Center) today sent the first commands to the HTV simulator at the PROX ground site via ISS.

The CDR completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Greg handled the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Significant Events Ahead (all dates Eastern Time, some changes possible.):
11/02/08 — Progress 30P reboost; Daylight Saving Time (DST) ends
11/14/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/16/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/25/08 — Progress M-65/30P undocking & deorbit
11/26/08 — Progress M-66/31P launch
11/29/08 — STS-126/Endeavour/ULF2 landing (~2:00 pm EST est.)
11/30/08 — Progress M-66/31P docking
02/09/09 — Progress M-66/31P undocking & deorbit
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
02/14/09 — STS-119/Discovery/15A docking
02/24/09 — STS-119/Discovery/15A undocking
02/26/09 — STS-119/Discovery/15A landing (nominal)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 — Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress M-67/32P undocking & deorbit
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/25/09 — Soyuz TMA-15/19S launch
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking)
07/30/09 — STS-128/Atlantis/17A – MPLM (P), last crew rotation
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P)
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

SpaceRef staff editor.